# Structure of the Cayley-Dickson algebras

### v1

@@ 5 Split Cayley-Dickson Algebras
$u_{2}$ $u_{4}$ $u_{14},u_{34},u_{1234}$ $u_{13}$ $u_{14}$ $-u_{124}$
$u_{2}$ $u_{14}$ $-u_{134},-u_{234}$ $u_{23}$ $u_{4}$ $u_{14}$
$u_{2}$ $u_{34}$ $u_{134}$ $u_{23}$ $u_{24}$ $-u_{124}$
Split-sedenions, ${\rm\widetilde{\mathbb{U}}_{1}}$, have 180 zero divisor pairs derived from 108 X, 48 A and 24 C-non-associative triads which includes extending the zero divisor to allow for unitary products. Unlike the split octonions, ${\rm\widetilde{\mathbb{U}}_{1}}$ has a full cyclic structure with correspondence of the $d$ terms for all cycle triads. The X types have only two modes and not necessarily the dual. The A and C types either have only the primary mode or all four modes with A type having three such cases and C have just two. This reduces to the 39 primaries shown in Table 15 and with 21 modes then the 3 triad cycles for each triad gives the 180 zero divisor pairs for ${\rm\widetilde{\mathbb{U}}_{1}}$. The $d$ blades shown as negative in the table require $(-a+b)(c+d)$ for all first cycles but some second and third cycles need the extended dual mode as well. The two exceptions are $(u_{1},u_{2},-u_{134})$ and $(u_{2},u_{3},-u_{124})$ which need the extended dual mode with the standard positive zero divisor.
An overlap between split and non-split zero divisors does not occur. The triad $(u_{12},u_{13},u_{14})$ and its product, $u_{12}u_{13}u_{14}=-u_{1234}$, contains terms that are all imaginary, which should be analogous to the non-split triad $(o_{12},o_{13},o_{14})$ but this has product $+o_{1234}$ providing a zero divisor with opposite parity.
@@ 6 Summary
The Cayley-Dickson algebras have been assumed to be symmetric in the permutation group but this is only valid for hypercomplex numbers. The octonions introduce non-associativity but in a symmetric form. Using a graded notation provides a rigorous definition to the order of basis elements and exposes the asymmetry of the power-associative Cayley-Dickson algebras. The octonion non-associative triads are alternate, which means swapping basis elements in the associator changes its sign. The term ultracomplex numbers has been introduced to cover the case of the associator toggling between zero and non-zero when swapping elements of the associator. This extends the symmetric non-associativity to another three types that exhibit asymmetric properties in cycles that leads to multiples of 84 zero divisor pairs and modes that, along with cycles, can reduce these down to 7, in some cases. The total number of zero divisors for each algebra is also derived.
The graded approach also applies to split ultracomplex numbers by showing that, like the hypercomplex case there is only one split algebra at each level of the Cayley-Dickson construction. In this case it is convenient to use $\epsilon=-1$ for all generators which then have positive square $u_{i}^{2}=1$. The convenient symmetry thereafter is that all odd length grades are unitary and all even grades are imaginary. The non-associativity from the non-split case carries over for identical indices with identical zero divisors for the imaginary terms whereas the unitary zero divisors can be related to ideals of the algebra or imaginary zero divisors. Open questions for future work are to establish connections of split zero divisors to ideals, gain further understand how zero divisors affect Malcev algebras, investigate the relationships between the pseudo-octonion algebras and establish why only three of the six are contained in power-associative algebras.
The author gratefully acknowledges the invaluable assistance and encouragement of James Chapell, Jacqui Ramagge and Derek Abbott for improving terminology, mathematical understanding and ongoing assistance with the publishing of this article. The author also acknowledges the support of an Australian Government Research Training Program Scholarship. All of this work was verified with the use of an octonion/ultronion calculator written in Python. The github URL for the calculator is

### v2

@@ 5 Split Cayley-Dickson Algebras
$u_{2}$ $u_{4}$ $u_{14},u_{34},u_{1234}$ $u_{13}$ $u_{14}$ $u_{124}$
$u_{2}$ $u_{14}$ $u_{134},-u_{234}$ $u_{23}$ $u_{4}$ $u_{14}$
$u_{2}$ $u_{34}$ $u_{134}$ $u_{23}$ $u_{24}$ $u_{124}$
Split sedenions, $\rm\widetilde{\mathbb{U}}_{1}$}, have 180 zero divisor pairs derived from 108 X, 48 A and 24 C-non-associative triads which includes extending the zero divisor to allow for unitary products. Unlike the split octonions, $\rm\widetilde{\mathbb{U}}_{1}$} has a full cyclic structure with correspondence of the $d$ terms for all cycle triads. The X types have only two modes and not necessarily the dual. The A and C types either have only the primary mode or all four modes with A type having three such cases and C have just two. This reduces to the 39 primaries shown in Table 15 and with 21 modes then the 3 triad cycles for each triad gives the 180 zero divisor pairs for $\rm\widetilde{\mathbb{U}}_{1}$}. The $d$ blades shown as negative in the table require (a+b)(c+d)(-a+b)(c+d) for all first cycles but some second and third cycles need the extended dual mode as well. The two exceptions are $(u_{1},u_{2},-u_{134})$ and $(u_{2},u_{3},-u_{124})$ which need the extended dual mode with the standard positive zero divisor.
An overlap between split and non-split zero divisors does not occur. The triad $(u_{12},u_{13},u_{14})$ and its product, $u_{12}u_{13}u_{14}=-u_{1234}$, contains terms that are all imaginary, which should be analogous to the non-split triad $(o_{12},o_{13},o_{14})$ but this has product $o_{1234}$ providing a zero divisor with opposite parity.
@@ 6 Summary
The Cayley-Dickson algebras have been assumed to be symmetric in the permutation of basis elements but this is only valid for hypercomplex numbers. The octonions introduce non-associativity but in a symmetric form. Using the graded notation provides a rigorous definition to the order of basis elements and exposes the asymmetry of the power-associative Cayley-Dickson algebras. The octonion non-associative triads are alternate, which means swapping basis elements in the associator changes its sign. The term ultracomplex numbers has been introduced to cover the case of the associator toggling between zero and non-zero when swapping elements of the associator. This extends the symmetric non-associativity to another three types that exhibit asymmetric properties in cycles that lead to multiples of 84 zero divisor pairs and modes that, along with cycles, can reduce these down to 7, if overlap is taken into account. The total number of zero divisors for each algebra is also derived.
The graded approach also applies to split ultracomplex numbers by showing that, like the hypercomplex case, there is only one split algebra at each level of the Cayley-Dickson construction. In this case it is convenient to use $\epsilon=-1$ for all generators, which then have positive square generators, $u_{i}^{2}=1$. The convenient symmetry thereafter is that all odd length grades are unitary and all even grades are imaginary. The non-associativity from the non-split case carries over for identical indices with identical zero divisors for the imaginary terms whereas the unitary zero divisors can be related to ideals of the algebra or imaginary zero divisors. Open questions for future work are to establish connections of split zero divisors to ideals, gain further understanding of how zero divisors affect Malcev algebras, investigate the relationships between the pseudo-octonion algebras and establish why only three of the six are contained in power-associative algebras.
The author gratefully acknowledges the invaluable assistance and encouragement of James Chapell, Jacqui Ramagge and Derek Abbott for improving terminology, mathematical understanding and ongoing assistance with the publishing of this article. Acknowledgement is also due to Maher ben abdessalem, Espirit, Tunisia for communicating a correction to $Z_{m}$. The author also acknowledges the support of an Australian Government Research Training Program Scholarship. All of this work was verified with the use of an octonion/ultronion calculator written in Python. The github URL for the calculator is